# Random regular graph states are complex at almost any depth

### v1

@@ 2.4 Drift and minorization of Markov chains
A function $V:\mathcal{X}\to[0,\infty]$ is a drift function (towards a set $C\subset\mathcal{X}$) for a Markov chain $P$ if there exists a constant $10<\lambda<$ 1 and a constant $b<\infty$ such that
@@
If a Markov chain has a drift function and satisfies the minorization condition on a small set $C$ in a compatible way, we can bound the convergence of the Markov chain to its stationary distribution in an exponential way. We will use the following theorem due to Rosenthal [58, Thm.12]:
@@
Let $P$ be a Markov chain on a (countable) state space $\mathcal{X}$, compatibly satisfying a drift and minorization condition with function $V(x)$, small set $C$ and parameters $\lambda,b,\delta,d$. Let $\nu$ be an initial distribution and let $\pi$ be the stationary distribution of $P$. We now have for all $10<r<$ 1 that:
@@ 3 Anticoncentration of graph states of constant degree
First, we compute the average second moment for a random multi-graph in the random pairing model. Conditioning on simplicity will then give results for uniformly random $d$-regular (simple) graphs, as discussed in Section 2.2. To obtain the result in LABEL:thm:anticoncentration for the random matching model, we will just need to slightly adapt the proof of this more complicated case.
@@
Consider the uniform measure $\mathcal{G}_{p}(n,d)$ on $d$-regular multi-graphs chosen from the pairing model. Then, for any $2<d=o(n)$ we have
@@
Corollary 18 (Restatement of Corollary 2).
@@
The corollary follows directly from Theorem 17, observing that the probability that a random multi-graph in the pairing model $G\leftarrow\mathcal{G}_{p}(d,n)$ is simple, is lower bounded by $1/2^{d^{2}}$, and hence, conditioning on the event that a graph in $\mathcal{G}_{p}(d,n)$ is simple yields the statement. ∎
@@ 3.1 Anticoncentration of a fixed graph state
The proof of Theorem 17 is based on a series of lemmata. As a first step, we find an expression for the second moment of a fixed graph states, averaged over the random choice of measurement angles. This follows more or less directly from the discussion in Appendix E of [36].
@@
of a ensemble of random states |$\ket{\psi}$ determines the second moment of its outcome distribution $p_{\psi}$ as $\mathbb{E}_{\psi}p_{\psi}(x)^{2}=\bra{x}^{\otimes 2}M_{2}\ket{x}^{\otimes 2}$, and thus computing the average second moment can be reduced to computing the second moment operator of the underlying state ensemble. For a random product rotation around the $Z$ axis applied to the |$\ket{+}^{\otimes n}$ state the second moments given by
$\mathbb{E}_{\theta}\left[\mathrm{e}^{-i\sum_{j}\theta_{j}Z_{j}}\ket{+^{n}}\bra{+^{n}}\mathrm{e}^{i\sum_{j}\theta_{j}Z_{j}}\right]=\frac{1}{4^{n}}({\sf I}+{\sf S}+{\sf X})^{\otimes n}=\frac{1}{4^{n}}\sum_{Q\in\mathcal{S}^{\otimes n}}Q,$ (25)
where ${\sf I}=\ket{01}\bra{01}+\ket{10}\bra{10}$ |, ${\sf S}=\ket{01}\bra{10}+\ket{10}\bra{01}$ |, ${\sf X}=\ket{00}\bra{00}+\ket{11}\bra{11}$ | and we let $\mathcal{S}=\{{\sf I},{\sf S},{\sf X}\}$ be the set of possible ‘states’ of the statmech model. Next, we observe that
@@
and that $\sum_{x\in\{0,1\}}\bra{xx}H^{\otimes 2}QH^{\otimes 2}\ket{xx}=1,\,\forall Q={\sf I},{\sf S},{\sf X}.$ Hence, the sign of a particular ‘state’ $Q\in\mathcal{S}^{\otimes n}$ on a graph |$\ket{G}$ is given by the parity of the $CZ$ gates that act on an ${\sf I}{\sf S}$ or ${\sf S}{\sf I}$ pair.
Thus, the second moment of the outcome distribution $p_{G,\theta}$ of measuring a graph state |$\ket{G}$ in the $X$-$Y$ plane angles $\theta$, averaged over the random choice of $\theta$, can be written as
@@
Importantly, this expression can also be interpreted for multi-graphs, where the elements of the adjacency matrix are now integers. The crucial number $|A_{G}[L,R]|$ featuring in Equation 23 thus still counts the number of edges crossing between the sets $L$ and $R$.
@@ 3.2 Averaging over matchings
Next, we need to evaluate the average over the random choice of (multi-)graphs. Since our graphs are defined in terms of uniformly random matchings, we begin by analyzing the average parity of edges in a random matching crossing between two disjoint sets $L,R\subset[n]$. We first note some symmetry properties of the average parity in Lemma 20, and then express the average parity in terms of Krawtchouk polynonmials in Lemma 21. In the following, let $\mathcal{M}(n)$ denotes the uniform distribution over perfect matchings of $n$ vertices.
@@
Consider a random matching $M$ on $n$ vertices and three disjoint sets $L,R,T\subseteq[n]$ covering $[n]=R\cup L\cup T$. The value of the average parity of the matching between two sets is invariant under the interchange of $R,L,T$ up to factors of ±$\pm 1$ as
@@
To see Equation 29, consider a particular matching that has $q$ edges going from $L$ to $R$ (and thus contributes $(-1)^{q}$ to the average). This matching also has $0\leq q^{\prime}\leq n-q$ edges going from $L$ to $T$. By construction the remaining vertices in $L$ must be matched to each other. Hence $q+q^{\prime}=|L|\mod 2$ and therefore $(-1)^{q}=(-1)^{q^{\prime}}(-1)^{|L|}$. Since this is true for every matching the claim follows. ∎
@@
Denote $T=[n]\backslash(R\cup L)$. Consider a matching that has $q$ edges between the sets $L$ and $R$ and $q^{\prime}$ edges between $L$ and $T$. Note that this is only possible if $q+q^{\prime}=|L|\mod 2$ and $q+q^{\prime}\leq|L|$. There are $\binom{|L|}{q}\binom{|R|}{q}\cdot q!$ ways to choose the first $q$ edges and $\binom{|L|-q}{q^{\prime}}\binom{|T|}{q^{\prime}}\cdot q^{\prime}!$ ways to choose the second set of edges. Once these edges are fixed there are $(|L|-q-q^{\prime}-1)!!(n-|L|-q-q^{\prime}-1)!!$! ways to complete this edge set to a full matching of $[n]$. Hence we can write
$\mathbb{E}_{M}(-1)^{|A_{M}(L,R)|}=\frac{1}{(n-1)!!}\sum_{\begin{subarray}{c}q,q^{\prime}\geq 0\\ q+q^{\prime}\leq|L|\\ q+q^{\prime}=|L|\mod 2\\ q\leq|R|\end{subarray}}(-1)^{q}q!\binom{|L|}{q}\binom{|R|}{q}q^{\prime}!\binom{|L|-q}{q^{\prime}}\binom{|T|}{q^{\prime}}\\ \times(|L|-q-q^{\prime}-1)!!(n-|L|-q-q^{\prime}-1)!!,$ (32)
since there are $(n-1)!!$! matchings in total. Since $\binom{|R|}{q}=0$ if $|R|<q$ we can drop the $q\leq|R|$ constraint going forward. We can expand the binomials involving $|L|$ and use the identity $(a-1)!!\,a!!=a!$ to rewrite this as
@@
3.3 Proof of Theorem 17
We can now provide a proof of the core result of this section (Theorem 17). This proof is long, but much of the conceptual has been done in the previous subsections (with ‘merely’ some counting remaining).
Proof of Theorem 17.
Lemmas 20 and 21 imply that all that is relevant for the average parity of the edges between two subsets $L,R\subseteq[n]$ is the size of the subsets. Recall that these sets correspond to ${\sf I}$ and ${\sf S}$ states on a subset of the vertices and let us think of the different states on the vertices as one of three colors ($L\leftrightarrow{\sf I}$, $R\leftrightarrow{\sf S}$, $T\leftrightarrow{\sf X}$). To analyze the random pairing model, we observe the following: for a fixed choice of $L,R\subseteq[n]$ in the sum (23), the average parity $\mathbb{E}_{{G_{M}}}[(-1)^{|A_{G_{M}}[L,R]|}]$ over graphs $G_{M}$ induced by a matching $M$ (the number of edges between ${\sf I}$-colored vertices ${\sf S}$-colored vertices) is equal to the average parity $\mathbb{E}_{M}[(-1)^{|A_{M}[L_{d},R_{d}]}]$ of two sets $L_{d},R_{d}\subseteq[dn]$ of size $|L_{d}|=|L|d,|R_{d}|=|R|d$ over the inducing matching. This is because we can ‘split up’ every vertex in $G_{M}$ into $d$ vertices of the same color. All that matters for the average parity is the crossings between colors blowing up the sets $L\rightarrow L_{d},R\rightarrow R_{d}$.
Hence, it follows from Lemmas 19, 20 and 21 that
@@ 3.3 Proof of Theorem 17
Let us denote (counter-intuitively, we are following the notation in [36]) $|L|=k,|R|=l$ (correspondingly $|L_{d}|=dk,|R_{d}|=dl$). Then we can use Lemma 21 to get
$\mathbb{E}$ [${}_{G_{M}\sim\mathcal{G}_{p}(n,d),\theta}\left[m_{2}(G_{M},\theta)\right]=2^{-n}\sum_{\begin{subarray}{c}L,R\subset[n]\\ L\cap R=\emptyset\end{subarray}}\sum_{\begin{subarray}{c}i=0\\ i=dk\mod 2\end{subarray}}^{dk}K^{d(n-k)}_{i}(dl)\frac{(dk)!(d(n-k)-i-1)!!}{(dk-i)!!(dn-1)!!}$ (37)
@@
where the second equality counts the number of possible choices of $L,R\subseteq[n]$ with fixed sizes $k,l$ and in the last equality we have implicitly defined $M_{p}(n,k,l,d)$. To get a feeling for what we need to show, consider Figure 2. There, we show the term $M_{p}(n,k,l,d)$ for fixed $n,d$ as a function of $k,l$. There is a clear separation of what is happening. At the boundaries, i.e., for $k=0$ or $l=0$ or $k=l$ the terms are largest, and they decay towards the center.
@@
Now, observe that $|R_{d}|=ld$ and therefore $(-1)^{ld}=1$ for even $d$ and $(-1)^{ld}=(-1)^{l}$ for odd $d$. By Lemma 20 $\mathbb{E}_{G_{M}\sim\mathcal{G}_{p}(n,d)}(-1)^{|A_{G_{M}}[\emptyset,R]|}=\mathbb{E}_{G_{M}\sim\mathcal{G}_{p}(n,d)}(-1)^{|A_{G_{M}}[\emptyset,[n]\setminus R]|}$ the second term thus vanishes for odd $d$ and is equal to the first term for even $d$.
We can now simplify the remaining sum to $|L|\leq n-|L|$ using Lemma 20 incurring a factor of $2$ and counting the number of choices of $L\subset n$
Case ${\mathrm{Case}\;0}=2^{-n}\sum_{l\leq n-l}\binom{n}{l}\mathbb{E}_{M\sim\mathcal{M}(nd)}(-1)^{|A_{M}[\emptyset,R_{d}]|}\cdot\begin{cases}4&\text{if }d=1\mod 2\\ 6&\text{ if }d=0\mod 2.\end{cases}$ (41)
Furthermore, we can use Lemma 21to get
@@
and the fact that $K^{dn}_{0}(dl)=1$ (see Equation 14) to get
Case ${\mathrm{Case}\;0}=2^{-n}\sum_{0\leq l\leq n}\binom{n}{l}\cdot\begin{cases}4/2&\text{ if }d=1\mod 2\\ 6/2&\text{ if }d=0\mod 2.\end{cases}\\ =\begin{cases}2&\text{ if }d=1\mod 2\\ 3&\text{ if }d=0\mod 2.\end{cases}$ (43)
@@
[…]
This set captures the terms in the sum where we need to achieve a nontrivial bound. This is because the sum over $k,l$ runs over at most $n^{2}$ terms and the average over the parity (the remaining part of Equation 46) is bounded by $1$ so that the sum over all $(k,l)\in W_{n}^{*}\setminus\mathcal{S}$ is bounded by $O(1/n)$. Using Lemma 20, we can always assume that $k\leq l\leq n-k-l$ (this gives a factor of $3!=6$ in the expression) and thus compute the remaining $k\geq 1$ terms of Equation 37 as
@@
For the first case, we will use the upper bound for Krawtchouck polynomials due to Derksen, given in Equation 17, in order to bound $M_{p}(n,k,l,d)$
@@
To further bound this, we consider the factors of Equation 49 individually
@@
[…]
Since $k^{d+1}\cdot 90^{d+1}\log(n)^{d+1}\leq n^{d/2}$ for $d>2$ and sufficiently large $n$, the terms in the sum are strictly decreasing in $k$. Hence, the sum is dominated by the $k=1$ term and we get for any $2<d=o(n)$
[…]
@@
$(n-1)!!$! $=\frac{n!}{2^{n/2}(n/2)!}\quad\Rightarrow\quad\frac{\sqrt{n!}}{(n-1)!!}=2^{n/2}\binom{n}{n/2}^{-1/2}$ (67)
@@
where we used that $(dk-i)!!\geq\sqrt{(dk-i)!}$ and that $h(1/2)=1$. Now we observe that this term is maximized at $i=\lfloor dk/2\rfloor$, since $i!(dk-i)!$ is minimized at $i=\lfloor dk/2\rfloor$. We thus get
@@
We can insert this back into Equation 46 and get
[…]
$\leq 6\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}\binom{n}{k}\binom{n-k}{l}\left[\sum_{\begin{subarray}{c}i=0\\ i=dk\mod 2\end{subarray}}^{dk}|K^{d(n-k)}_{i}(dl)|\frac{(dk)!(d(n-k)-i-1)!!}{(dk-i)!!(dn-1)!!}\right]$ (75)
$\leq\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}6dk2^{d(n-k)/2}\binom{n}{k}\binom{n-k}{l}\binom{d(n-k)}{dl}^{-\frac{1}{2}}\frac{(dk)!}{\lfloor dk/2\rfloor!}\frac{((d(n-k))!)^{\frac{1}{2}}}{(dn-1)!!}$ (76)
$\leq\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}6d^{3/2}k\sqrt{n}2^{d(n-k)/2}\binom{n}{k}\binom{n-k}{l}\binom{d(n-k)}{dl}^{-\frac{1}{2}}\binom{dn}{dk}^{-\frac{1}{2}}$ (77)
$\leq\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}Cd^{3/2}k\sqrt{n}2^{dn/2}\binom{n}{k}\binom{n-k}{l}\binom{d(n-k)}{dl}^{-\frac{1}{2}}\binom{dn}{dk}^{-\frac{1}{2}}$ (78)
$\leq\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}Cd^{3/2}k\sqrt{n}2^{dn/2}\binom{n}{k}\binom{n-k}{l}\binom{(n-k)}{l}^{-\frac{d}{2}}\binom{n}{k}^{-\frac{d}{2}}$ (79)
@@
This case is more subtle and corresponds to the dominant terms of the sum over $k,l$ which lie in a ‘ring’ of intermediate values, see Fig. 2(right). The first thing we note is that within this region there is a maximal value of $k$, achieved when $l=k$ and
@@
This equation is manifestly monotonously decreasing in $l$, since $l\leq(n-k)/2$ for any pair $(k,l)$. Hence we can insert $l_{*}$ to obtain an upper bound on Eq. 85. We further split the sum over $i$ into two cases: (i) $i\geq 130\log(n)$ and (ii) $i\leq 130\log(n)$. For case (i) we can use Hölder’s inequality
@@
We can upper bound the second factor by $1$ by noting that $\binom{d(n-k)}{i}=K_{i}^{d(n-k)}(0)$ and adding back in the $i\leq 130\log(n)$ terms to the sum. All of these terms are nonnegative since the only nonzero term at $x=0$ in the definition (14) of the Krawtchouk polynomials is the $q=0$ term. This allows us to recognize the formula for the average parity over all matchings from Lemma 21 with $l=0$ which trivially evaluates to $1$. For the maximization we note that $0\leq i/d\leq k\leq k^{*}=c^{*}n$ and $l_{*}\geq c_{*}n$, telling us that
@@
where we have used that $1/\binom{dn}{dk}$ is monotonously decreasing in $k$, so that we can bound the binomial using lower bound $k_{*}=90\log(n)$, and the last bound follows from the assumption $2<d=o(n)$. Hence we have that
[…]
@@
Putting the three bounds together, we have our result that for $2<d\in o(n)$

### v2

@@ 2.4 Drift and minorization of Markov chains
A function $V:\mathcal{X}\to[0,\infty]$ is a drift function (towards a set $C\subset\mathcal{X}$) for a Markov chain $P$ if there exists a constant $0<\lambda<1$ and a constant $b<\infty$ such that
@@
If a Markov chain has a drift function and satisfies the minorization condition on a small set $C$ in a compatible way, we can bound the convergence of the Markov chain to its stationary distribution in an exponential way. We will use the following theorem due to Rosenthal [78, Thm.12]:
@@
Let $P$ be a Markov chain on a (countable) state space $\mathcal{X}$, compatibly satisfying a drift and minorization condition with function $V(x)$, small set $C$ and parameters $\lambda,b,\delta,d$. Let $\nu$ be an initial distribution and let $\pi$ be the stationary distribution of $P$. We now have for all $0<r<1$ that:
@@ 3 Anticoncentration of graph states of constant degree
First, we compute the average second moment for a random multi-graph in the random pairing model. Conditioning on simplicity will then give results for uniformly random $d$-regular (simple) graphs, as discussed in Section˜2.2. To obtain the result in LABEL:thm:anticoncentration for the random matching model, we will just need to slightly adapt the proof of this more complicated case.
@@
Consider the uniform measure $\mathcal{G}_{p}(n,d)$ on $d$-regular multi-graphs chosen from the pairing model. Then, for any $2<d=o(n^{1/2})$ we have
@@
Corollary 18 (Restatement of Corollary˜2).
@@
The corollary follows directly from Theorem˜17, observing that the probability that a random multi-graph in the pairing model $G\leftarrow\mathcal{G}_{p}(d,n)$ is simple, is lower bounded by $1/2^{d^{2}}$, and hence, conditioning on the event that a graph in $\mathcal{G}_{p}(d,n)$ is simple yields the statement. ∎
@@ 3.1 Anticoncentration of a fixed graph state
The proof of Theorem˜17 is based on a series of lemmata. As a first step, we find an expression for the second moment of a fixed graph states, averaged over the random choice of measurement angles. This follows more or less directly from the discussion in Appendix E of [51].
@@
[…]
@@
[…]
@@
Importantly, this expression can also be interpreted for multi-graphs, where the elements of the adjacency matrix are now integers. The crucial number $|A_{G}[L,R]|$ featuring in Equation˜23 thus still counts the number of edges crossing between the sets $L$ and $R$.
@@ 3.2 Averaging over matchings
Next, we need to evaluate the average over the random choice of (multi-)graphs. Since our graphs are defined in terms of uniformly random matchings, we begin by analyzing the average parity of edges in a random matching crossing between two disjoint sets $L,R\subset[n]$. We first note some symmetry properties of the average parity in Lemma˜20, and then express the average parity in terms of Krawtchouk polynonmials in Lemma˜21. In the following, let $\mathcal{M}(n)$ denotes the uniform distribution over perfect matchings of $n$ vertices.
@@
Consider a random matching $M$ on $n$ vertices and three disjoint sets $L,R,T\subseteq[n]$ covering $n]=R\cup L\cup T$. The value of the average parity of the matching between two sets is invariant under the interchange of $R,L,T$ up to factors of $\pm 1$ as
@@
To see Equation˜29, consider a particular matching that has $q$ edges going from $L$ to $R$ (and thus contributes $(-1)^{q}$ to the average). This matching also has $0\leq q^{\prime}\leq n-q$ edges going from $L$ to $T$. By construction the remaining vertices in $L$ must be matched to each other. Hence $q+q^{\prime}=|L|\mod 2$ and therefore $(-1)^{q}=(-1)^{q^{\prime}}(-1)^{|L|}$. Since this is true for every matching the claim follows. ∎
@@
Denote $T=[n]\backslash(R\cup L)$. Consider a matching that has $q$ edges between the sets $L$ and $R$ and $q^{\prime}$ edges between $L$ and $T$. Note that this is only possible if $q+q^{\prime}=|L|\mod 2$ and $q+q^{\prime}\leq|L|$. There are $\binom{|L|}{q}\binom{|R|}{q}\cdot q!$ ways to choose the first $q$ edges and $\binom{|L|-q}{q^{\prime}}\binom{|T|}{q^{\prime}}\cdot q^{\prime}!$ ways to choose the second set of edges. Once these edges are fixed there are $(|L|-q-q^{\prime}-1)!!(n-|L|-q-q^{\prime}-1)!!$ ways to complete this edge set to a full matching of $[n]$. Hence we can write
$\mathbb{E}_{M}(-1)^{|A_{M}(L,R)|}=\frac{1}{(n-1)!!}\hskip-20.00003pt\sum_{\begin{subarray}{c}q,q^{\prime}\geq 0\\ q+q^{\prime}\leq|L|\\ q+q^{\prime}=|L|\mod 2\\ q\leq|R|\end{subarray}}\hskip-20.00003pt(-1)^{q}q!\binom{|L|}{q}\binom{|R|}{q}q^{\prime}!\binom{|L|-q}{q^{\prime}}\binom{|T|}{q^{\prime}}\\ \times(|L|-q-q^{\prime}-1)!!(n-|L|-q-q^{\prime}-1)!!,$ (32)
since there are (n1)!!(n-1)!! matchings in total. Since $\binom{|R|}{q}=0$ if $|R|<q$ we can drop the $q\leq|R|$ constraint going forward. We can expand the binomials involving $|L|$ and use the identity (a1)!!a!!=a!(a-1)!!\,a!!=a! to rewrite this as
@@
3.3 Proof of Theorem˜17
We can now provide a proof of the core result of this section (Theorem˜17). This proof is long, but much of the conceptual has been done in the previous subsections (with ‘merely’ some counting remaining).
Proof of Theorem˜17.
Lemmas˜20 and 21 imply that all that is relevant for the average parity of the edges between two subsets $L,R\subseteq[n]$ is the size of the subsets. Recall that these sets correspond to $\sf I$} and $\sf S$} states on a subset of the vertices and let us think of the different states on the vertices as one of three colors ($L\leftrightarrow{\sf I}$, $R\leftrightarrow{\sf S}$, $T\leftrightarrow{\sf X}$). To analyze the random pairing model, we observe the following: for a fixed choice of $L,R\subseteq[n]$ in the sum (23), the average parity $\mathbb{E}_{{G_{M}}}[(-1)^{|A_{G_{M}}[L,R]|}]$ over graphs $G_{M}$ induced by a matching $M$ (the number of edges between $\sf I$}-colored vertices $\sf S$}-colored vertices) is equal to the average parity $\mathbb{E}_{M}[(-1)^{|A_{M}[L_{d},R_{d}]}]$ of two sets $L_{d},R_{d}\subseteq[dn]$ of size $|L_{d}|=|L|d,|R_{d}|=|R|d$ over the inducing matching. This is because we can ‘split up’ every vertex in $G_{M}$ into $d$ vertices of the same color. All that matters for the average parity is the crossings between colors blowing up the sets $L\rightarrow L_{d},R\rightarrow R_{d}$.
Hence, it follows from Lemmas˜19, 20 and 21 that
@@ 3.3 Proof of Theorem˜17
Let us denote (counter-intuitively, we are following the notation in [51]) $|L|=k,|R|=l$ (correspondingly $|L_{d}|=dk,|R_{d}|=dl$). Then we can use Lemma˜21 to get
$\mathbb{E}$ $_{G_{M}\sim\mathcal{G}_{p}(n,d),\theta}\left[m_{2}(G_{M},\theta)\right]=2^{-n}\sum_{\begin{subarray}{c}L,R\subset[n]\\ L\cap R=\emptyset\end{subarray}}\sum_{\begin{subarray}{c}i=0\\ i=dk\mod 2\end{subarray}}^{dk}K^{d(n-k)}_{i}(dl)\frac{(dk)!(d(n-k)-i-1)!!}{(dk-i)!!(dn-1)!!}$ (37)
@@
where the second equality counts the number of possible choices of $L,R\subseteq[n]$ with fixed sizes $k,l$ and in the last equality we have implicitly defined $M_{p}(n,k,l,d)$. To get a feeling for what we need to show, consider Figure˜2. There, we show the term $M_{p}(n,k,l,d)$ for fixed $n,d$ as a function of $k,l$. There is a clear separation of what is happening. At the boundaries, i.e., for $k=0$ or $l=0$ or $k=l$ the terms are largest, and they decay towards the center.
@@
Now, observe that $|R_{d}|=ld$ and therefore $(-1)^{ld}=1$ for even $d$ and $(-1)^{ld}=(-1)^{l}$ for odd $d$. By Lemma˜20 $\mathbb{E}_{G_{M}\sim\mathcal{G}_{p}(n,d)}(-1)^{|A_{G_{M}}[\emptyset,R]|}=\mathbb{E}_{G_{M}\sim\mathcal{G}_{p}(n,d)}(-1)^{|A_{G_{M}}[\emptyset,[n]\setminus R]|}$ the second term thus vanishes for odd $d$ and is equal to the first term for even $d$.
We can now simplify the remaining sum to $|L|\leq n-|L|$ using Lemma˜20 incurring a factor of $2$ and counting the number of choices of $L\subset n$
${\mathrm{Case}\;0}=2^{-n}\sum_{l\leq n-l}\binom{n}{l}\mathbb{E}_{M\sim\mathcal{M}(nd)}(-1)^{|A_{M}[\emptyset,R_{d}]|}\cdot\begin{cases}4&\text{ if }d=1\mod 2\\ 6&\text{ if }d=0\mod 2.\end{cases}$ (41)
Furthermore, we can use Lemma˜21to get
@@
and the fact that $K^{dn}_{0}(dl)=1$ (see Equation˜14) to get
[…]
@@
$\mathcal{S}$ $\coloneqq\bigg\{(k,l)\in W^{*}_{n}\;\bigg|\;\;2^{-n}\binom{n}{k}\binom{n-k}{l}\geq n^{-3}\bigg\}.$ (45)
This set captures the terms in the sum where we need to achieve a nontrivial bound. This is because the sum over $k,l$ runs over at most $n^{2}$ terms and the average over the parity (the remaining part of Equation˜46) is bounded by $1$ so that the sum over all $(k,l)\in W_{n}^{*}\setminus\mathcal{S}$ is bounded by $O(1/n)$. Using Lemma˜20, we can always assume that $k\leq l\leq n-k-l$ (this gives a factor of $3!=6$ in the expression) and thus compute the remaining $k\geq 1$ terms of Equation˜37 as
@@
For the first case, we will use the upper bound for Krawtchouck polynomials due to Derksen, given in Equation˜17, in order to bound $M_{p}(n,k,l,d)$
@@
To further bound this, we consider the factors of Equation˜49 individually
@@
$2^{n}|T_{\mathrm{Case}\;1}|\leq 6\sum_{k=1}^{90\log(n)}2^{n-k}\frac{n^{k+1/2}}{\sqrt{n-k}}dk(dk)!C_{4}^{dk}\log(n)^{dk+1}n^{-dk/2}\\ \leq 6\sum_{k=1}^{90\log(n)}2^{n-k}\frac{n^{k+1/2}}{\sqrt{n-k}}(C_{4})^{dk}\log(n)^{dk+1}(dk)^{dk+1}n^{-dk/2}.$ (63)
Since $k^{d+1}\cdot 90^{d+1}\log(n)^{d+1}\leq n^{d/2}$ for $d>2$ and sufficiently large $n$, the terms in the sum are strictly decreasing in $k$. Hence, the sum is dominated by the $k=1$ term and we get for any $2<d=o(n^{1/2})$
$|T_{\mathrm{Case}\;1}|\in O\left(d^{d+1}\log(n)^{d+3}n^{-d/2+1}\right)=o(1),$ (64)
@@
$(n-1)!!$ $=\frac{n!}{2^{n/2}(n/2)!}\quad\Rightarrow\quad\frac{\sqrt{n!}}{(n-1)!!}=2^{n/2}\binom{n}{n/2}^{-1/2}$ (67)
@@
where we used that $(dk-i)!!\geq\sqrt{(dk-i)!}$ and that $h(1/2)=1$. Now we observe that this term is maximized at $i=\lfloor dk/2\rfloor$, since i!(dki)!i!(dk-i)! is minimized at $i=\lfloor dk/2\rfloor$. We thus get
@@
We can insert this back into Equation˜46 and get
$2^{n}|$ $T_{\mathrm{Case}\;2}|$
$\leq 6\hskip-21.33955pt\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}\binom{n}{k}\binom{n-k}{l}\left[\sum_{\begin{subarray}{c}i=0\\ i=dk\mod 2\end{subarray}}^{dk}|K^{d(n-k)}_{i}(dl)|\frac{(dk)!(d(n-k)-i-1)!!}{(dk-i)!!(dn-1)!!}\right]$ (75)
$\leq\hskip-21.33955pt\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}\hskip-18.00005pt6dk2^{d(n-k)/2}\binom{n}{k}\binom{n-k}{l}\binom{d(n-k)}{dl}^{-\frac{1}{2}}\frac{(dk)!}{\lfloor dk/2\rfloor!}\frac{((d(n-k))!)^{\frac{1}{2}}}{(dn-1)!!}$ (76)
$\leq\hskip-21.33955pt\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}\hskip-18.00005pt6d^{3/2}k\sqrt{n}2^{d(n-k)/2}\binom{n}{k}\binom{n-k}{l}\binom{d(n-k)}{dl}^{-\frac{1}{2}}\binom{dn}{dk}^{-\frac{1}{2}}$ (77)
$\leq\hskip-21.33955pt\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}\hskip-18.00005ptCd^{3/2}k\sqrt{n}2^{dn/2}\binom{n}{k}\binom{n-k}{l}\binom{d(n-k)}{dl}^{-\frac{1}{2}}\binom{dn}{dk}^{-\frac{1}{2}}$ (78)
$\leq\hskip-21.33955pt\sum_{\begin{subarray}{c}(k,l)\in\mathcal{S},\\ 8\log(n)\leq k\leq l\leq n-k-l\end{subarray}}\hskip-18.00005ptCd^{3/2}k\sqrt{n}2^{dn/2}\binom{n}{k}\binom{n-k}{l}\binom{(n-k)}{l}^{-\frac{d}{2}}\binom{n}{k}^{-\frac{d}{2}}$ (79)
@@
This case is more subtle and corresponds to the dominant terms of the sum over $k,l$ which lie in a ‘ring’ of intermediate values, see Fig.˜2(right). The first thing we note is that within this region there is a maximal value of $k$, achieved when $l=k$ and
@@
This equation is manifestly monotonously decreasing in $l$, since $l\leq(n-k)/2$ for any pair $(k,l)$. Hence we can insert $l_{*}$ to obtain an upper bound on Eq.˜85. We further split the sum over $i$ into two cases: (i) $i\geq 130\log(n)$ and (ii) $i\leq 130\log(n)$. For case (i) we can use Hölder’s inequality
@@
We can upper bound the second factor by $1$ by noting that $\binom{d(n-k)}{i}=K_{i}^{d(n-k)}(0)$ and adding back in the $i\leq 130\log(n)$ terms to the sum. All of these terms are nonnegative since the only nonzero term at $x=0$ in the definition (14) of the Krawtchouk polynomials is the $q=0$ term. This allows us to recognize the formula for the average parity over all matchings from Lemma˜21 with $l=0$ which trivially evaluates to $1$. For the maximization we note that $0\leq i/d\leq k\leq k^{*}=c^{*}n$ and $l_{*}\geq c_{*}n$, telling us that
@@
where we have used that $1/\binom{dn}{dk}$ is monotonously decreasing in $k$, so that we can bound the binomial using lower bound $k_{*}=90\log(n)$, and the last bound follows from the assumption $2<d=o(n^{1/2})$. Hence we have that
$|T_{\mathrm{Case}\;3}|\leq n^{2}n^{8}\cdot O(n^{-10.5}+n^{-\Omega(\log n)})\leq O(n^{-1/2})=o(1).$ (102)
@@
Putting the three bounds together, we have our result that for $2<d\in o(n^{1/2})$